In a water tank there is 12.0 liters of water at a temperature of 44.0 °C. Boiling water at 100.0 °C is poured into the tank until the water is at 90.0 °C. How many liters of boiling water was poured into the tank? The specific heat capacity of liquid water is 4.19 kJ/kg/°C.

1
Expert's answer
2022-02-09T16:17:50-0500

volume of water in the tank = "12l =0.012m^3"

density of water "=1000kg\/m^3"

mass of water in the tank = density "\times" volume "=0.012\times 1000=12kg"

let the water added be "x kg"

heat lost by water at 100"^0C=" heat gained by water up to "90^0C"

"x \times 4.19\times 1000 (100-90)=(12)\times 4.19\times 1000(90-44)"

"10x=12\times 46=552"

"x=55.2kg"

volume of hot water added ="mass\/density=55.2\/1000=0.0552m^3=55.2litres"